# Bridging the Culture Gap: A Framework for LLM-Driven Socio-Cultural Localization of Math Word Problems in Low-Resource Languages

### v1

@@ 2 Related Work
Although LLMs are highly capable of solving complex tasks, they often struggle with simple variations in input (askari2025assessing; chhabra2024revisiting; achara2025watching). For example, Abedin et al. (2025) demonstrate that even minor perturbations such as spelling mistakes can significantly impact model performance, a challenge that is particularly pronounced in low-resource settings and evaluations.
@@ 3.1 Pipeline
LLMs for Localization. In this work, we leveraged Gemini models Comanici et al. (2025) due to their exceptional multilingual performance shown by natural language generation (NLG) multilingual benchmarks Ojo et al. (2023). To reduce experimental costs while assessing our approach, we used Gemini-1.5-pro for evaluation, and employed Gemini-2.5-pro for the final data generation.
@@ 3.2 Datasets Used
To ensure translation quality, we filtered the outputs using SSA-COMMET Li et al. (2025), a sentence-level semantic similarity metric that scores translation quality on a scale from 0 to 1, with 1 indicating perfect translation. Based on the score distribution shown in the Appendix C, we retained only translations with a COMMET score above 0.65. From these, we selected the top 1,500 examples per language for automatic localization and use as training data in our after localizing each. While intermediate experiments leverage Gemini-1.5-pro to evaluate the pipeline and assess its effectiveness due to cost constraints and to avoid unwanted behaviors, such as prompt caching [^4] https://ai.google.dev/gemini-api/docs/caching?lang=python, the final dataset is generated using Gemini-2.5-pro to ensure the highest quality standards.
@@ 5 Results and Analysis
Additionally, we find that languages with more training resources tend to benefit more from localization. This trend is evident from the table, where languages are sorted in ascending order based on their availability in the FineWeb2 Penedo et al. (2025) dataset. Higher-resource languages (appearing lower in the table) show greater performance gains, while lower-resource languages (toward the top) experience more limited improvements. More exploration to see characteristics of languages affected by the finetuning experiment included in the Appendix B.
@@ 6 Conclusion
Appendices
@@
Localization means adapting the text to the target culture and context by: Identifying and replacing English-specific entities (such as person names, organization names, and currency symbols/names) with culturally appropriate equivalents in {target_lang_name}. Ensuring the localized text sounds natural and grammatically correct in {target_lang_name}. Crucially, the localized text should be generated directly from the English sentence, without relying on any pre-existing native translation.
English Sentence to Localize: {original_eng}
Your Goal: Produce a Localized {target_lang_name} sentence.
Instructions: Identify person names, organization names, and currency symbols/names in the English sentence. Replace these entities with culturally appropriate {target_lang_name} alternatives. If an entity does not have a direct cultural equivalent or is already culturally neutral, it may remain unchanged. Respond with ONLY the Localized {target_lang_name} sentence and nothing else.
Localized {target_lang_name}:
Appendix A More Results
@@
Appendix B Training Dataset Statstics
@@
Appendix C Translation quality - SSA-COMMET score

### v2

@@ 2 Related Work
Although LLMs are highly capable of solving complex tasks, they often struggle with simple variations in input (Askari et al., 2025; Chhabra et al., 2024; Achara and Chhabra, 2025). For example, Abedin et al. (2025) demonstrate that even minor perturbations such as spelling mistakes can significantly impact model performance, a challenge that is particularly pronounced in low-resource settings and evaluations.
@@ 3.1 Pipeline
LLMs for Localization. In this work, we leveraged Gemini models Comanici et al. (2025) due to their exceptional multilingual performance shown by natural language generation (NLG) multilingual benchmarks Ojo et al. (2023). To reduce experimental costs while assessing our approach, we used Gemini-1.5-pro for evaluation of the framework, and employed Gemini-2.5-pro for the final data generation.
@@ 3.2 Datasets Used
To ensure translation quality, we filtered the outputs using SSA-COMMET Li et al. (2025), a sentence-level semantic similarity metric that scores translation quality on a scale from 0 to 1, with 1 indicating perfect translation. Based on the score distribution shown in the Appendix D, we retained only translations with a COMMET score above 0.65. From these, we selected the top 1,500 examples per language for automatic localization and use as training data in our after localizing each. While intermediate experiments leverage Gemini-1.5-pro to evaluate the pipeline and assess its effectiveness due to cost constraints and to avoid unwanted behaviors, such as prompt caching [^4] https://ai.google.dev/gemini-api/docs/caching?lang=python, the final dataset is generated using Gemini-2.5-pro to ensure the highest quality standards.
@@ 5 Results and Analysis
Additionally, we find that languages with more training resources tend to benefit more from localization. This trend is evident from the table, where languages are sorted in ascending order based on their availability in the FineWeb2 Penedo et al. (2025) dataset. Higher-resource languages (appearing lower in the table) show greater performance gains, while lower-resource languages (toward the top) experience more limited improvements. More exploration to see characteristics of languages affected by the finetuning experiment included in the Appendix C.
@@ 6 Conclusion
Acknowledgment
The authors would like to thank the German Federal Ministry of Education and Research and the German federal states (http://www.nhr-verein.de/en/our-partners) for supporting this work/project as part of the National High-Performance Computing (NHR) joint funding program. We also thank PaliGemma academic program for gemini credits used in this research.
Appendix
Appendix A Prompt Templates Used
Below, we present the prompts used for our Auto-Localizer and direct prompt localization. We observed that including the intermediate steps shown in Table 1 improves the accuracy of localization.
@@ Appendix A Prompt Templates Used
Localization means adapting the text to the target culture and context by: Identifying and replacing English-specific entities (such as person names, organization names, and currency symbols/names) with culturally appropriate equivalents in {target_lang_name}. Ensuring the localized text sounds natural and grammatically correct in {target_lang_name}. Crucially, the localized text should be generated directly from the English sentence, without relying on any pre-existing native translation. English Sentence to Localize: {original_eng} Your Goal: Produce a Localized {target_lang_name} sentence. Instructions: Identify person names, organization names, and currency symbols/names in the English sentence. Replace these entities with culturally appropriate {target_lang_name} alternatives. If an entity does not have a direct cultural equivalent or is already culturally neutral, it may remain unchanged. Respond with ONLY the Localized {target_lang_name} sentence and nothing else. Localized {target_lang_name}:
Appendix B Additional Evaluation Results
Figure 6 presents additional evaluations, extending the analysis to a broader set of models with diverse characteristics.
@@ Appendix B Additional Evaluation Results
Appendix C Precence of Cultural entities in our Training data
Figure 7 illustrates the distribution of culturally relevant items within the 1,500 data samples selected for each language. Our analysis reveals that a considerable proportion of the translated items could not be localized to the target language, primarily because they did not contain the types of entities, such as personal names, organizations, or currencies, that form the basis of our localization strategy. This highlights a key limitation of translation-only benchmarks: when source items lack culturally grounded entities, they fail to provide meaningful opportunities for localization, thereby reducing the effectiveness of culturally adapted datasets. By quantifying this gap, we emphasize the importance of designing datasets that explicitly incorporate entity-level cultural relevance to ensure robust evaluation and training across languages.
@@ Appendix C Precence of Cultural entities in our Training data
Appendix D Translation quality - SSA-COMMET score
@@
Appendix E Code and Reproducibility
We used ElutherAI’s open source Language Model Evaluation Harness (lm-eval) framework Gao et al. (2024) to evaluate models. Instead of direct model querying this allows us to have standardized reproducible evaluations across all LLMs. We also set generation parameters (i.e. temperature) to zero for consistency.